OpenStudy (anonymous):

can someone help me with the steps factoring -6y^2+53y+9

OpenStudy (anonymous):

When factored it'll look like this: -(y-9) (6 y+1)

OpenStudy (anonymous):

ok can you help me with the steps please

OpenStudy (ybarrap):

Using, http://upload.wikimedia.org/wikipedia/commons/thumb/3/3d/Completing_the_square.ogv/400px--Completing_the_square.ogv.jpg with a=-6, b=53 and c = 9 a(y + b/2a)^2 + c - (b/2a)^2 = -6y^2+53y+9 -6(y +53/2(-6))^2 + 9 - (53/2*(-6))^2 That's it!

OpenStudy (anonymous):

Okay. So first you'd look for something that can be used with (-6) and 9. So it'd be: -6y^2 - 54y +1y +9 =(-6y^2 - 54y) + (1y + 9) Common factor from each bracket and it results in: =-6y(y+9) +1 (y+9) =(-6y + 1) (y+9)

OpenStudy (anonymous):

ok thank you